OpenStudy (anonymous):

factor this: x^8-98x^4+1 I dont want just the answer I want to see steps to how to get the answer

OpenStudy (anonymous):

if you make b = x^4 what is the equation you'd have ?

OpenStudy (anonymous):

b^2 - 98b + 1 so whay do you do next?

OpenStudy (anonymous):

it becomes a quadratic

OpenStudy (anonymous):

plug it into the quadratic formula or complete the square

OpenStudy (anonymous):

okay when you complete the square you get this (b-49)^2 -2400 what do you do next?

OpenStudy (anonymous):

i think that once you solve for b, plug x^4 back in

OpenStudy (anonymous):

sorry i want it to be perfectly factored in the form (x+a)(x+b)

OpenStudy (anonymous):

hmmmm

OpenStudy (anonymous):

srry man im lost...im still learning this stuff too..i was trying to figure it out while you were

OpenStudy (anonymous):

yea its fine thanks for helping

OpenStudy (anonymous):

um...use the general factor form. \[(x^2 - \frac{ -b + \sqrt{b^2 - 4ac} }{ 2a })(x^2 - \frac{ -b - \sqrt{b^2 - 4ac} }{ 2a })\] where a=1, b=-98, c=1. voila!

OpenStudy (anonymous):

oh yeah forgot about that lol

OpenStudy (anonymous):

Though it's quite ugly, it's reliable. lol Good luck! :)